if tan A = 2/3 and if angle A is acute, then sin A = ?
Answer by stanbon(48558)

if tan A = 2/3 and if angle A is acute, then sin A = ?
-----------------
tan A = y/x; y = 2; x = 3
Then r = sqrt(2^2+3^2) = sqrt(13)
sin A = y/x = 2/sqrt(13)
=====================
Cheers,
Stan H.